OpenStudy (pandaecat):

An exponential growth function never exceeds a quadratic function with a positive leading coefficient. True or false?

OpenStudy (pandaecat):

@mathmate @Nnesha @phi

OpenStudy (phi):

do you know any example of an exponential function?

OpenStudy (pandaecat):

Yes, like g(x)=x^2

OpenStudy (phi):

do you know the example of a quadratic function?

OpenStudy (pandaecat):

h(x)=x^2+4x+12

OpenStudy (phi):

g(x)=x^2 is also quadratic exponential is for example 2^x

OpenStudy (pandaecat):

Or 3^x? Oh right I got it messed up

OpenStudy (phi):

for easy computation let's use g(x)= 10^x for the exponential and f(x)= x^2 for the quadratic. when x= 3, 10^3 = 1000 and x^2 = 3*3= 9 is the exponential bigger than the quadratic when x=3 ? answer: yes

OpenStudy (phi):

so **An exponential growth function never exceeds a quadratic function?** is not true. i.e FALSE

OpenStudy (pandaecat):

Oh ty!